# Time-dependent dynamics in the confined lattice Lorentz gas

### v1

@@ I Introduction
Experiments on active-microrheology are nowadays largely employed in order to infer properties of soft materials [1, 2, 3]. In these experiments, a tracer particle is pulled through a medium by means of either optical or magnetic tweezers. The application of this experimental procedure finds applications in several contexts ranging from colloidal suspensions [3], soft glassy materials [4, 5], fluid interfaces [6], to living cells [7]. The possibility to tune the external force allows for the exploration of the full crossover from the linear regime to the truly non-linear one emerging for strong forces and its associated non-equilibrium phenomena such as force-thinning [8, 9, 10] and enhanced diffusivities [11, 12, 13]. In addition to these phenomena, among the most fascinating challenges for the theory is the description of effects due to biasing, crowding, and confinement [14, 15] in strongly interacting systems out of equilibrium. Several analytical progresses have been obtained within theoretical models in which a tracer particle moves in a dense environment [16, 17, 18]; see [19, 20, 21, 22, 23, 24, 25] for exact results for lattice models. Despite these achievements, an analytically tractable model able to encompass the out-of-equilibrium regime of strong pulling in the presence of crowing and geometrical confinement is still missing to date.
@@
A first attempt in this direction has been put forward by the driven quasi-confined lattice Lorentz gas model whose analytic solution is presented in the accompanying Letter [26]. This model describes the complementary situation in which a tracer particle explores an immobile landscape of randomly distributed obstacles. This minimal model of disorder is motivated by the fact that the mutual exclusion between the tracer and the environment is commonly regarded as the most important ingredient. The confinement in a lattice model can be realized in several ways: the simplest is referred to as quasi-confinement, which is implemented by restricting the model onto an infinitely long strip with identified edges; this choice corresponds to the topology of a cylinder of finite circumference $L$. For this lattice model we shall elaborate an analytical solution, valid to first order in the obstacle density and for arbitrarily strong driving and confinement size $L$. The theory encompasses the full crossover from the maximally confined $L=2$ lane model to the unbounded model ($L=\infty$) whose solution in equilibrium is known for a long time by the pioneering works of Nieuwenhuizen et al. in Refs. [27, 28, 29, 30, 31, 32, 33], while the non-equilibrium case has been solved more recently [34, 35, 36, 37, 38]. The analytical method, developed in [34, 35, 36], and here applied to the quasi-confined model, relies on a lattice version of the scattering formalism borrowed from quantum mechanics [39].
@@
The scope of this paper is to derive and discuss the complete time-dependent dynamics, the approach to the non-equilibrium stationary state, and how these aspects interplay with the presence of a confinement. The principal results of this paper can be summarized as follows: the stationary state is characterized by a terminal velocity which is approached – for infinitesimally small forces – as a power-law $t^{-d_{E}/2}$ whose exponent is $d_{E}$ for any finite value of $L$, while $d_{E}=2$ for the unbounded model [34]. The above long-time tails are actually fragile since an arbitrarily small but finite force is sufficient to decorated them by an exponential decay. The existence of an effective dimensionality $d_{E}$ is actually the signature of a dimensional crossover. This surprising behavior, that we find already at equilibrium, can be characterized by the complete time-dependent velocity autocorrelation function (VACF). The latter exhibits a long-time tail of the form $t^{-(d_{E}+2)/2}$ [33] with effective dimension $d_{E}=2$ at short times, while for large times the effective dimension $d_{E}=1$ entails a reduction to a quasi one-dimensional system. We then study the fluctuations of the tracer particle along the force. From the variance of the position we infer the emergence of a transient superdiffusive behavior at intermediate times with a time-dependent exponent of anomalous diffusion that can saturates to $\alpha(t)=3$ for large forces. Standard diffusion, i.e., $\alpha(t)=1$, is attained at both short times and long times. At long times however the diffusion is characterized by a force-dependent and confinement-dependent diffusion coefficient that we identify. Among our findings is that, in general, for large forces the confinement effects fade out and the system behaves, de facto, as the unbounded one. On the other hand, confinement-induced effects play a major role at low forces. Their effect is so strong that it can alter the non-analytic behavior of response functions making them significantly different from the one observed in the unbounded model. This is what we observed by studying the long-time diffusion coefficient.
@@
This paper is organized as follows. In Sec. II we introduce the driven, quasi-confined, lattice Lorentz gas model. The exact solution of the model relies on the scattering approach formalism elaborated in [34, 35, 36] that we briefly recall in Sec. III. The results for the velocity response are presented in Sec. IV. Sec. V conveys the analysis of the VACF in equilibrium and its associated dimensional crossover, as well as the time-dependent fluctuations along the force. Our summary and conclusions are provided in Sec. VI. A series of Appendices collects several technicalities including Green’s function of the quasi-confined model.
@@ II The model
The lattice Lorentz gas model consists in a random walk occurring on a disordered lattice $\Lambda$ in which some sites are not accessible to the walker. Let $\Lambda=\{\mathbf{r}=(x,y):x=1,\ldots,L_{x},y=1,\ldots,L_{y}\}$ with $N=L_{x}L_{y}$ sites and periodic boundary conditions. The walker performs jumps to its nearest neighbors, $\mathcal{N}=\{\pm\textbf{e}_{x},\pm\textbf{e}_{y}\}$. The waiting time at accessible sites is Poisson-distributed with mean waiting time $\tau$. The disorder is modeled via hard and immobile obstacles (also referred to as impurities). If the tracer attempts to hop on a site occupied by an obstacle, then the transition is rejected and the walker remains where it was (still time is progressing and the walker doesn’t move); the trajectory is schematically depicted in Fig. 1. Throughout this paper we will consider periodic boundary conditions along both the lattice directions. Technically speaking, the limit $L_{x}\rightarrow\infty$ is anticipated while the size $L$ will be kept finite. This procedure defines the notion of quasi-confinement adopted in this paper.
@@
The state of the system at time $t$ is completely characterized by the site occupation probability density $p_{\textbf{r}}(t)$ which expresses the probability to be at site r at time $t$. We formally introduce a Hilbert space $\mathcal{H}$ with orthonormal basis states of position kets $\{|\textbf{r}\rangle:\textbf{r}\in\Lambda\}$. We then define an abstract ket state $|p(t)\rangle\in\mathcal{H}$ by $|p(t)\rangle=\sum_{\textbf{r}\in\Lambda}p_{\textbf{r}}(t)|\textbf{r}\rangle$. Using completeness we find
@@
[…]
$\partial_{t}\langle\textbf{r}|p(t)\rangle=\sum_{\textbf{r}\in\Lambda}\langle\textbf{r}|\hat{H}|\textbf{r}^{\prime}\rangle|\textbf{r}^{\prime}\rangle\langle\textbf{r}|p(t)\rangle\,,$ (4)
@@
[…]
@@ III.1 Bare dynamics
By adopting the bra-ket notation the conditional probability is written as $U_{0}(\textbf{r},t|\textbf{r}^{\prime})=\langle\textbf{r}|\hat{U}_{0}(t)|\textbf{r}^{\prime}\rangle$ in the complete orthonormal basis of position kets with time-evolution operator $\hat{U}_{0}(t)$. The master equation for the time-evolution operator reads $\partial_{t}\hat{U}_{0}(t)=\hat{H}_{0}\hat{U}_{0}(t)$, with initial condition $𝟙\hat{U}_{0}(t=0)=\openone$ and $\hat{H}_{0}$ given by Eq. (5). The time-evolution operator formally solves the master equation by means of $\hat{U}_{0}(t)=\exp(\hat{H}_{0}t)$. The site-occupation probability at time $t$ given that the driving force $F>0$ is switched on at time $t=0$ is given by
@@
$\epsilon_{x}(k)=\frac{1}{2}(1-\cosh(F/2))-\frac{1}{2}(1-\cos(k+\mathsf{i}F/2))\,.$ (15)
@@
We note that as long as $L_{x}$ is finite the momenta $k_{x}$ form a discrete set but since the limit $L_{x}\rightarrow\infty$ is anticipated it follows that such momenta form a continuous; this allows us to perform the momentum derivatives $\partial_{k_{x}}$ as shown in Eq. (20). Parenthetically, we observe that $D_{x}^{0}(F)$ does not depend on the value of the size $L$.
@@ III.2 Obstructed biased motion
We consider the presence of hard, immobile obstacles randomly distributed on the lattice. The effects due the disordered lattice can be worked out exactly to first order in the obstacle density $n$ and for an arbitrary strength $F$ of the bias. In this section we we review the analytical technique based on the application of the scattering formalism in quantum mechanics (see e.g. [39]) and originally developed in [34, 41] for the lattice Lorentz gas.
When the tracer attempts to jump onto an obstacle site, the tracer remains at its initial site before the jump. Thus, the obstacle cancels the transition from and to the unaccessible site. In order to take into account the obstacles, we split the Hamiltonian into the sum of a free contribution $\hat{H}_{0}$ describing the dynamics on the empty lattice and an interaction term $\hat{V}$ accounting for the disorder, i.e., $\hat{H}=\hat{H}_{0}+\hat{V}$. Let $\hat{v}(\textbf{s}_{i})\equiv\hat{v}_{i}$ be the potential for a single impurity located at site $\textbf{s}_{i}$. Let $N_{I}$ be the number of such obstacles on the lattice, then the total potential $\hat{V}$ can be written as a sum over all single-obstacle potentials $\hat{v}_{i}$ with $\hat{V}=\sum_{i=1}^{N_{I}}\hat{v}_{i}$, the density of the obstacle is thus $n=N_{I}/N$. Technically, a tracer starting its motion on an empty site will remain on an empty site, while if the motion would have started on a cluster of touching obstacles, then the trajectory will remain in such a cluster.
Let us consider now an isolated obstacle that we put in site $\textbf{s}_{1}$. However, since the system is translationally invariant, we may pick any site as origin, so we do for $\textbf{s}_{1}$. Since the tracer performs nearest-neighbor jumps, the only non-vanishing matrix elements for an obstacle located in the origin are the obstacle site $\textbf{e}_{0}=\textbf{0}$ and its nearest neighbors $\textbf{e}_{\pm 2}=\pm\textbf{e}_{y}$, $\textbf{e}_{\pm 1}=\pm\textbf{e}_{x}$. Cancellation of the transition rates entails that matrix elements $\langle\textbf{e}_{\mu}|\hat{v}|\textbf{e}_{\nu}\rangle$ of the single-obstacle potential are $\langle\textbf{0}|\hat{v}|\textbf{0}-\textbf{d}\rangle=-\Gamma W(\textbf{d})$ and $\langle\textbf{0}-\textbf{d}|\hat{v}|\textbf{0}\rangle=-\Gamma W(-\textbf{d})$ where $\textbf{d}\in\mathcal{N}$, where $\mathcal{N}$ denotes the neighboring sites of the obstacle (see the discussion at the end of Sec.III). The matrix representation of the single obstacle potential in the basis $\{\textbf{e}_{-2},\textbf{e}_{-1},\textbf{e}_{0},\textbf{e}_{1},\textbf{e}_{2}\}$ reads
@@
The full propagator $\hat{G}(s)=(s-\hat{H})^{-1}$ and the free propagator $\hat{G}_{0}(s)=(E-\hat{H}_{0})^{-1}$, written as resolvents, are related via
@@
The first sum in Eq. (26) takes into account the repeated scatterings with a single impurity, this term contributes in first order of the density $n$. The remaining sums appearing in the multiple scattering expansion [Eq. (26)] correspond to contributions involving correlated scattering with at least two impurities, these terms contribute to order $O(n^{2})$ and higher. Throughout this paper we will not be interested in the effects due to a specific arrangement of the lattice impurities, therefore we will perform an average over the disorder realizations, this is denoted [.$[.]_{\rm av}$. The averaging procedure restores the translational invariance and, as a result, the disorder-averaged scattering operator $\langle\textbf{k}|[\hat{T}]_{\rm av}|\textbf{k}\rangle$ is diagonal in the plane-wave basis. The first term in the scattering expansion [Eq. (26)] is identified as the one yielding a linear contribution in the density $n$. The remaining terms are responsible for corrections of order $O(n^{2})$ or higher and are interpreted as effects arising from correlated scatterings between different impurities. Retaining the first sum in the scattering expansion amounts to the so-called single-scatterer approximation [42]. This approximation scheme yields $\langle\textbf{k}|[\hat{T}]_{\rm av}|\textbf{k}\rangle=nN\langle\textbf{k}|\hat{t}|\textbf{k}\rangle+O(n^{2})$, where $\langle\textbf{k}|\hat{t}|\textbf{k}\rangle:=t(\textbf{k})$ is the forward scattering amplitude of a single obstacle and $nN=N_{I}$ is the number of impurities. From Eq. (25) we obtain the disorder-average propagator in plane wave basis, $[G]_{\text{av}}(\textbf{k}):=\langle\textbf{k}|[\hat{G}]_{\text{av}}|\textbf{k}\rangle$, exact to first order in the obstacle density
@@
The first term in the above is the bare propagator in Fourier space and Laplace domain, therefore $G_{0}(\textbf{k})=\langle\textbf{k}|\hat{G}_{0}|\textbf{k}\rangle=\int_{0}^{\infty}\textrm{d}t\,e^{-st}\langle\textbf{k}|\hat{U}_{0}(t)|\textbf{k}\rangle=[s-\epsilon(\textbf{k})]^{-1}$, the latter is diagonal in the plane-wave basis. From now on the dependence on the Laplace parameter $s$ will be suppressed for notational convenience. The second term in Eq. (27) is the self-energy $\Sigma(\textbf{k})=nNt(\textbf{k})+O(n^{2})$ to linear order in the density $n$, the latter follows from Dyson’s equation, $[G_{\rm av}](\textbf{k})=\left(G_{0}(\textbf{k})^{-1}-\Sigma(\textbf{k})\right)^{-1}$.
@@
$[G]_{\text{av}}^{\text{c}}(\textbf{k})=G_{0}(\textbf{k})+n[G_{0}(\textbf{k})+Nt(\textbf{k})G_{0}(\textbf{k})^{2}]+O(n^{2})\,.$ (28)
The disorder-average propagator encodes all moments of the displacement of the tracer particle to first order in the density and for arbitrarily strong driving. Results at linear order in $n$ are due to single scatterings and are completely encoded in the solution of the scattering problem for a single impurity. The scattering operator $\hat{t}$ for a single obstacle satisfies

### v2

@@ I Introduction
Experiments on active-microrheology are nowadays largely employed in order to infer properties of soft materials [1, 2, 3]. In these experiments, a tracer particle is pulled through a medium by means of either optical or magnetic tweezers. The application of this experimental procedure finds applications in several contexts ranging from colloidal suspensions [3], soft glassy materials [4, 5], fluid interfaces [6], to living cells [7]. The possibility to tune the external force allows for the exploration of the full crossover from the linear regime to the truly non-linear one emerging for strong forces and its associated non-equilibrium phenomena such as force-thinning [8, 9, 10] and enhanced diffusivities [11, 12, 13]. In addition to these phenomena, among the most fascinating challenges for the theory is the description of effects due to biasing, crowding, and confinement [14, 15] in strongly interacting systems out of equilibrium.
Several analytical results have been obtained for lattice models in which a tracer particle moves in a dense environment [16, 17, 18]. In particular, for single-file diffusion there exist exact [19, 20] and numerical results [21, 22, 23, 24] as well as approximate [25] and experimental ones [26], including also the theoretical analysis of the biased dynamics [27] and the motion in an external potential [28]. Furthermore, analytical progresses have been made also on closely related fronts of the literature. This includes – inter alia – exclusion processes [29, 30, 31, 32] and their variants with a moving defect [33, 34], driven one-dimensional lattice gases [35, 36] in narrow channels [37, 38, 39, 40, 41], diffusion on percolation networks [42, 43, 44] [^1] See [69] for a rigorous proof of the results of [42]., and the random trapping for a tracer moving into a channel [46]. Despite these achievements, an analytically tractable model able to encompass the out-of-equilibrium regime of strong pulling in the presence of crowing, disorder, and geometrical confinement is still missing to date.
@@
This paper aims at providing first attempt in this direction through the driven quasi-confined lattice Lorentz gas model. This model describes the complementary situation in which a tracer particle explores an immobile landscape of randomly distributed obstacles. This minimal model of disorder is motivated by the fact that the mutual exclusion between the tracer and the environment is commonly regarded as the most important ingredient. The confinement in a lattice model can be realized in several ways: the simplest is referred to as quasi-confinement, which is implemented by restricting the model onto an infinitely long strip with identified edges; this choice corresponds to the topology of a cylinder of finite circumference $L$.
For this lattice model we shall elaborate an analytical solution, valid to first order in the obstacle density, $n$, and for arbitrarily strong driving and confinement size. In this regard, all formulas should be read including order $O(n)$. Strictly speaking, this approach cannot describe finite $n$, a regime that is far beyond the scope of the paper; hence, conclusions cannot be drawn for finite densities at infinite times. To this end, it is important to stress a further intricacy. The limits of $n\rightarrow 0$ and $F\rightarrow\infty$ (with $t\rightarrow\infty$) do not commute, this makes the scenario rather subtle. However, the simulations we performed show in a clear fashion the domain of applicability of the theory and, anticipating some results, we specify that the window of validity is not uniform as it depends on $F$ and $n$. We therefore compute the first-order correction for arbitrary force and then, for practical purposes, we compare it with simulations protracting to large although finite times, a circumstance that applies to practical scenarios. The theory encompasses the full crossover from the maximally confined $L=2$ lane model to the unbounded model ($L=\infty$) whose solution in equilibrium is known for a long time by the pioneering works of Nieuwenhuizen et al. in Refs. [47, 48, 49, 50, 51, 52, 53], while the non-equilibrium case has been solved more recently [54, 55, 56, 56, 57, 58]. The analytical method, developed in [54, 55, 56], and here applied to the quasi-confined model, relies on a lattice version of the scattering formalism borrowed from quantum mechanics [59].
@@
The scope of this paper is to derive and discuss the complete time-dependent dynamics, the approach to the non-equilibrium stationary state, and how these aspects interplay with the presence of a confinement. The principle results of this paper can be summarized as follows: the stationary state is characterized by a terminal velocity which is approached – for infinitesimally small forces – as the power-law $t^{-1/2}$ for any finite value of $L$, whereas the power law is $t^{-1}$ for the unbounded model [54]. The above long-time tails are actually fragile since an arbitrarily small but finite force is sufficient to decorate them by an exponential decay; i.e., for the confined system the long time-tail becomes $t^{-1/2}\exp(-cF^{2}t)$, where $c>0$ is a coefficient. The existence of an effective dimensionality is actually the signature of a dimensional crossover. This surprising behavior, that we find already at equilibrium, can be characterized by the complete time-dependent velocity autocorrelation function (VACF). The latter exhibits a long-time tail of the form $t^{-(d+2)/2}$ [53] with effective dimension $d=2$ at short times, while for large times the effective dimension $d=1$ entails a reduction to a quasi one-dimensional system. We then study the fluctuations of the tracer particle along the force. From the variance of the position we infer the emergence of a transient superdiffusive behavior at intermediate times with a time-dependent exponent of anomalous diffusion that can saturates to $\alpha(t)=3$ for large forces. Standard diffusion, i.e., $\alpha(t)=1$, is attained at both short times and long times. At long times however the diffusion is characterized by a force-dependent and confinement-dependent diffusion coefficient that we identify. Among our findings is that, in general, for large forces the confinement effects fade out and the system behaves, de facto, as the unbounded one. On the other hand, confinement-induced effects play a major role at low forces. Their effect is so strong that it can alter the non-analytic behavior of response functions making them significantly different from the one observed in the unbounded model. This is what we observed by studying the long-time diffusion coefficient.
@@
This paper is organized as follows. In Sec. II we introduce the driven, quasi-confined, lattice Lorentz gas model. The exact solution of the model relies on the scattering approach formalism elaborated in [54, 55, 56] that we briefly recall in Sec. III. The results for the velocity response are presented in Sec. IV. Section V conveys the analysis of the VACF in equilibrium and its associated dimensional crossover, as well as the time-dependent fluctuations along the force. Our summary and conclusions are provided in Sec. VI. A series of Appendices collects several technicalities including Green’s function of the quasi-confined model.
@@ II The model
The lattice Lorentz gas model is a random walk occurring on a disordered lattice $\Lambda$ in which some sites are not accessible to the walker. Let $\Lambda=\{\mathbf{r}=(x,y):x=1,\ldots,L_{x},y=1,\ldots,L_{y}\}$ with $N=L_{x}L_{y}$ sites and periodic boundary conditions. The walker performs jumps to its nearest neighbors, $\mathcal{N}=\{\pm\textbf{e}_{x},\pm\textbf{e}_{y}\}$. The waiting time at accessible sites is exponentially distributed with mean waiting time $\tau$. The disorder is modeled via hard and immobile obstacles (also referred to as impurities). If the tracer attempts to hop on a site occupied by an obstacle, then the transition is rejected and the walker remains where it was (still time is progressing and the walker doesn’t move); the trajectory is schematically depicted in Fig. 1. Throughout this paper we will consider periodic boundary conditions along both the lattice directions. Technically speaking, the limit $L_{x}\rightarrow\infty$ is anticipated while the size $L\coloneq L_{y}$ will be kept finite. This procedure defines the notion of quasi-confinement adopted in this paper.
@@
The state of the system at time $t$ is completely characterized by the site-occupation probability density $p_{\textbf{r}}(t)$ which expresses the probability to be at site r at time $t$. We formally introduce a Hilbert space $\mathcal{H}$ with orthonormal basis states of position kets $\{|\textbf{r}\rangle:\textbf{r}\in\Lambda\}$. We then define an abstract ket state $|p(t)\rangle\in\mathcal{H}$ by $|p(t)\rangle=\sum_{\textbf{r}\in\Lambda}p_{\textbf{r}}(t)|\textbf{r}\rangle$. Using completeness we find
@@
Hence, $p_{\textbf{r}}(t)=\langle\textbf{r}|p(t)\rangle$ is the probability to find the tracer at site r at time $t$. The time evolution of this probability follows according to the master equation $\partial_{t}|p(t)\rangle=\hat{H}|p(t)\rangle$. The latter is formally analogous to the Schrödinger equation in imaginary time provided $\hat{H}$ is identified as the “Hamiltonian”. Upon projecting the master equation onto the bra $\langle\textbf{r}|$ we find the following representation of the master equation in the position basis
$\partial_{t}\langle\textbf{r}|p(t)\rangle=\sum_{\textbf{r}\in\Lambda}\langle\textbf{r}|\hat{H}|\textbf{r}^{\prime}\rangle\langle\textbf{r}^{\prime}|p(t)\rangle\,$, (4)
@@
The choice of this prefactor corresponds to non-normalized transition rates considered in [55, 56], while normalizad rates have been considered in [54]. The two choices are actually equivalent under a rescaling of the time. One may choose units of time such that the mean waiting time $\tau=1$ sets the time scale. The presence of hard immobile obstacles precluding transitions from and to impurities can be modeled by adding an interaction potential that cancels exactly the associated transition rates. The Hamiltonian is formally written as $\hat{H}=\hat{H}_{0}+\hat{V}$, where the “potential” $\hat{V}$ is constructed such that its matrix elements cancel the non-allowed transitions. For a single obstacle located at site s the potential takes the form $\hat{V}=\hat{v}(\textbf{s})\equiv\hat{v}_{1}$. The nonvanishing matrix elements are those between the state associated to the site s and its neighbors: $\textbf{s}-\textbf{d}$, with $\textbf{d}\in\mathcal{N}$. For example, $\langle\textbf{s}|\hat{v}_{1}|\textbf{s}-\textbf{d}\rangle=-\Gamma W(\textbf{d})$ cancels the transition to the impurity, while $\langle\textbf{s}-\textbf{d}|\hat{v}_{1}|\textbf{s}\rangle=-\Gamma W(-\textbf{d})$ cancels the transition from the impurity, and $\langle\textbf{s}-\textbf{d}|\hat{v}_{1}|\textbf{s}-\textbf{d}\rangle=\Gamma W(\textbf{d})$.
@@ III.1 Bare dynamics
[…]
@@
$\epsilon_{x}(k)=\frac{1}{2}[1-\cosh(F/2)]-\frac{1}{2}[1-\cos(k+\mathsf{i}F/2)]\,.$ (15)
@@
We note that as long as $L_{x}$ is finite the momenta $k_{x}$ form a discrete set but since the limit $L_{x}\rightarrow\infty$ is anticipated it follows that such momenta form a continuum; this allows us to perform the momentum derivatives $\partial_{k_{x}}$ as shown in Eq. (20). Parenthetically, we observe that $D_{x}^{0}(F)$ does not depend on the value of the size $L$.
@@ III.2 Obstructed biased motion
We consider the presence of hard, immobile obstacles randomly distributed on the lattice. The effects due the disordered lattice can be worked out exactly to first order in the obstacle density $n$ and for an arbitrary strength $F$ of the bias. In this section we review the analytical technique based on the application of the scattering formalism in quantum mechanics (see e.g. [59]) and originally developed in [54, 56, 56] for the lattice Lorentz gas.
When the tracer attempts to jump onto an obstacle site, the tracer remains at its initial site before the jump. Thus, the obstacle cancels the transition from and to the unaccessible site. In order to take into account the obstacles, we split the Hamiltonian into the sum of a free contribution $\hat{H}_{0}$ describing the dynamics on the empty lattice and an interaction term $\hat{V}$ accounting for the disorder, i.e., $\hat{H}=\hat{H}_{0}+\hat{V}$. Let $\hat{v}(\textbf{s}_{i})\equiv\hat{v}_{i}$ be the potential for a single impurity located at site $\textbf{s}_{i}$. Let $N_{I}$ be the number of such obstacles on the lattice, then the total potential $\hat{V}$ can be written as a sum over all single-obstacle potentials $\hat{v}_{i}$ with $\hat{V}=\sum_{i=1}^{N_{I}}\hat{v}_{i}$, the density of the obstacle is thus $n=N_{I}/N$. Technically, a tracer beginning on an empty site will remain within the connected cluster of empty sites.
Let us consider now an isolated obstacle that we put in site $\textbf{s}_{1}$. However, since the system is translationally invariant, we may pick any site as origin, so we do for $\textbf{s}_{1}$. Since the tracer performs nearest-neighbor jumps, the only non-vanishing matrix elements for an obstacle located in the origin are the obstacle site $\textbf{e}_{0}=\textbf{0}$ and its nearest neighbors $\textbf{e}_{\pm 2}=\pm\textbf{e}_{y}$, $\textbf{e}_{\pm 1}=\pm\textbf{e}_{x}$. Cancellation of the transition rates entails that matrix elements $\langle\textbf{e}_{\mu}|\hat{v}|\textbf{e}_{\nu}\rangle$ of the single-obstacle potential are $\langle\textbf{0}|\hat{v}|\textbf{0}-\textbf{d}\rangle=-\Gamma W(\textbf{d})$ and $\langle\textbf{0}-\textbf{d}|\hat{v}|\textbf{0}\rangle=-\Gamma W(-\textbf{d})$ where $\textbf{d}\in\mathcal{N}$, where $\mathcal{N}$ denotes the neighboring sites of the obstacle (see the discussion at the end of Sec.III). The matrix representation of the single obstacle potential in the basis $(\textbf{e}_{-2},\textbf{e}_{-1},\textbf{e}_{0},\textbf{e}_{1},\textbf{e}_{2})$ reads
@@
The full propagator $\hat{G}(s)=(s-\hat{H})^{-1}$ and the free propagator $\hat{G}_{0}(s)=(s-\hat{H}_{0})^{-1}$, written as resolvents, are related via
@@
The first sum in Eq. (26) takes into account the revisit of a certain impurity, this term contributes in first order of the density $n$. The remaining sums appearing in the multiple scattering expansion [Eq. (26)] correspond to contributions involving correlated scattering with at least two impurities, these terms contribute to order $O(n^{2})$ and higher. Throughout this paper we will not be interested in the effects due to a specific arrangement of the lattice impurities, therefore we will perform an average over the disorder realizations, this is denoted $_{\rm av}$. The averaging procedure restores the translational invariance and, as a result, the disorder-averaged scattering operator $\hat{T}]_{\rm av}$ is diagonal in the plane-wave basis $|\textbf{k}\rangle$. The first term in the scattering expansion [Eq. (26)] is identified as the one yielding a linear contribution in the density $n$. The remaining terms are responsible for corrections of order $O(n^{2})$ or higher and are interpreted as effects arising from correlated scatterings between different impurities. Retaining the first sum in the scattering expansion amounts to the so-called single-scatterer approximation [61]. This approximation scheme yields $\langle\textbf{k}|[\hat{T}]_{\rm av}|\textbf{k}\rangle=nN\langle\textbf{k}|\hat{t}|\textbf{k}\rangle+O(n^{2})$, where $\langle\textbf{k}|\hat{t}|\textbf{k}\rangle:=t(\textbf{k})$ is the forward scattering amplitude of a single obstacle and $nN=N_{I}$ is the number of impurities. From Eq. (25) we obtain the disorder-average propagator in plane wave basis, $G]_{\text{av}}(\textbf{k}):=\langle\textbf{k}|[\hat{G}]_{\text{av}}|\textbf{k}\rangle$, exact to first order in the obstacle density
@@
The first term in the equation above is the bare propagator in Fourier space and Laplace domain, therefore $G_{0}(\textbf{k})=\langle\textbf{k}|\hat{G}_{0}|\textbf{k}\rangle=\int_{0}^{\infty}\textrm{d}t\,e^{-st}\langle\textbf{k}|\hat{U}_{0}(t)|\textbf{k}\rangle=[s-\epsilon(\textbf{k})]^{-1}$, the latter is diagonal in the plane-wave basis. From now on the dependence on the Laplace parameter $s$ will be suppressed for notational convenience. The second term in Eq. (27) is the self-energy $\Sigma(\textbf{k})=nNt(\textbf{k})+O(n^{2})$ to linear order in the density $n$, the latter follows from Dyson’s equation, $G_{\rm av}](\textbf{k})=\left[G_{0}(\textbf{k})^{-1}-\Sigma(\textbf{k})\right]^{-1}$.
@@
$G]_{\text{av}}^{\text{c}}(\textbf{k})=G_{0}(\textbf{k})+n[G_{0}(\textbf{k})+Nt(\textbf{k})G_{0}(\textbf{k})^{2}]+O(n^{2})\,.$ (28)
The disorder-averaged propagator encodes all moments of the tracer’s displacement to first order in the obstacle density $n$ and is valid for arbitrarily strong driving. At linear order in $n$, one must include sequences in which the particle undergoes arbitrarily many scatterings at a given obstacle, then moves to a different obstacle and scatters there arbitrarily many times, and so on – subject to the crucial constraint that no previously visited obstacle is revisited after the tracer has interacted with a different obstacle. Processes that violate this non-revisiting condition (i.e., returns to an earlier obstacle after encountering another) contribute only beyond first order in $n$ and are therefore excluded at this level. Consequently, the linear-in-$n$ scattering series is built from free propagations interspersed with single-obstacle scattering episodes, each of which is completely characterized by the single-impurity $t$-matrix. The single-obstacle scattering operator $\hat{t}$ satisfies